While conducting an independent-measures t-test, a researcher reports a t statistic with df = 18, how many individual subjects participated in the experiment?  

a. 20

b. 18

c. cannot be determined from the information given
Answer #1

Solution :

Given that,

Degrees of freedom = df =18

sample size = n = df + 1 = 18 + 1 = 19

c. cannot be determined from the information given